OpenStudy (anonymous):

Sorry to ask a stupid question but in the second lecture (Kinematics) there was this equation (x = 8 - 6t + t2) and i never studied this before, could some one explain how to solve this, Thank you

OpenStudy (anonymous):

x is the displacement. it is given as a function of time 't'. now velocity v=dx/dt differentiate the given function w.r.t time, you will get the velocity. if you need to find acceleration, differentiate it again.

OpenStudy (anonymous):

Do you know derive? If not tell me I teach you the simpliest way.

OpenStudy (sk8rgirl12):

It's not stupid. You don't know unless you ask or try. One of the 2. Either experience it, or learn from someone who has. That is my best advice.

OpenStudy (anonymous):

we have displacement as function of time \[x = 8 - 6t + t^2 \] and when you differentiate it, you get \[frac{dx}{dt} = -6 + 2t = v \] and when you differentiate it again you get \[frac{dv}{dt} = 2 = a \] a is acceleration and v is velocity, here you have simple derivatives like \[frac{d(x^2)}{dx} = 2x \] etc...

OpenStudy (anonymous):

sorry for my math formula bad writing, i am not familiar with this syntax (this is not latex ?) :) frac is just a fraction